# Infrared-Assisted Single-Stage Framework for Joint Restoration and Fusion of Visible and Infrared Images under Hazy Conditions

### v2

@@ IV-C Comparison with State-of-the-art Methods
Additionally, we conduct a quantitative comparison on 361 image pairs from the MSRS dataset to verify the effectiveness of our method. Tables I and II present the experimental results based on the DIACMP and Dehazeformer dehazing methods, respectively. As shown in Tables III, our method ranks first across all eight metrics, demonstrating its outstanding performance on the MSRS dataset. The higher $Q_{AB/F}$ and $Q_{CV}$ scores indicate that our fused images achieve superior detail clarity. Meanwhile, the best $Q_{SCD}$ and $Q_{VIF}$ scores suggest that our method ensures better visual consistency, closely matching the clear source images. The $Q_{MI}$ metric reflects the shared information between the fused and source images, confirming that our approach preserves more source image information. As perceptual quality metrics, $Q_{NIQE}$ and $Q_{PI}$ evaluate image naturalness and perceptual quality in a no-reference manner, where lower values indicate better alignment with human visual perception. Furthermore, $Q_{SF}$ measures the sharpness of image edges, demonstrating that our method produces haze-free fused images that are both natural and sharp.
@@
Figure 3: Visual comparison of fusion results from different methods on the MSRS dataset. In the fusion results generated by the comparison methods, except for the last column, the first row of each image pair represents the result of first dehazing with DIACMP and then fusing. The second row shows the result of dehazing with Dehazeformer and then fusing. The last column represents the result of fusion using Text IF (first row) and our method (second row).
@@
TABLE I: Quantitative analysis of our method compared with diacmp+fusion and text-if on the MSRS dataset. Red indicates the best, and blue indicates the second best.
@@
Methods $Q_{MI}$$\uparrow$ $Q_{AB/F}$$\uparrow$ $Q_{CV}$$\downarrow$ $Q_{SCD}$$\uparrow$ $Q_{VIF}$$\uparrow$ $Q_{NIQE}$$\downarrow$ $Q_{PI}$$\downarrow$ $Q_{SF}$$\uparrow$ MLFusion 1.706 0.274 608.391 1.121 0.199 5.022 4.285 9.953 U2Fusion 1.301 0.330 838.330 1.364 0.226 4.977 4.354 7.209 LRRNet 1.864 0.441 614.101 1.000 0.288 4.750 4.333 8.479 ALFusion 1.449 0.377 700.180 1.201 0.257 5.391 5.419 7.417 TIMFusion 1.800 0.382 1048.770 1.090 0.295 4.515 4.256 8.136 MRFS 1.580 0.457 332.068 1.348 0.325 5.068 4.800 9.042 Text-IF 1.428 0.558 458.252 1.351 0.369 4.260 4.312 8.393 Ours 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050
@@
TABLE II: Quantitative analysis of our method compared with dehazeformer+fusion and text-if on the MSRS dataset. Red indicates the best, and blue indicates the second best.
@@
Methods $Q_{MI}$$\uparrow$ $Q_{AB/F}$$\uparrow$ $Q_{CV}$$\downarrow$ $Q_{SCD}$$\uparrow$ $Q_{VIF}$$\uparrow$ $Q_{NIQE}$$\downarrow$ $Q_{PI}$$\downarrow$ $Q_{SF}$$\uparrow$ MLFusion 1.720 0.270 613.291 1.119 0.199 5.052 4.337 9.826 U2Fusion 1.311 0.324 832.878 1.361 0.225 5.185 4.439 7.071 LRRNet 1.883 0.436 608.019 0.999 0.288 4.883 4.445 8.317 ALFusion 1.471 0.370 695.816 1.198 0.256 5.468 5.542 7.339 TIMFusion 1.785 0.382 1057.931 1.092 0.299 4.567 4.355 8.303 MRFS 1.603 0.448 328.901 1.340 0.323 5.150 4.905 8.843 Text-IF 1.428 0.558 458.252 1.351 0.369 4.260 4.312 8.393 Ours 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050
@@
Figure 4: Visual comparison of fusion results from different methods on the M3FD dataset. In the fusion results generated by the comparison methods, except for the last column, the first row of each image pair represents the result of first dehazing with DIACMP and then fusing. The second row shows the result of dehazing with Dehazeformer and then fusing. The last column represents the result of fusion using Text IF (first row) and our method (second row).
@@
TABLE III: Quantitative analysis of our method compared with diacmp+fusion and text-if on the M3FD dataset. Red indicates the best, and blue indicates the second best.
@@
Methods $Q_{MI}$$\uparrow$ $Q_{AB/F}$$\uparrow$ $Q_{CV}$$\downarrow$ $Q_{SCD}$$\uparrow$ $Q_{VIF}$$\uparrow$ $Q_{NIQE}$$\downarrow$ $Q_{PI}$$\downarrow$ $Q_{SF}$$\uparrow$ MLFusion 1.842 0.411 937.069 1.134 0.390 4.840 3.230 12.025 U2Fusion 1.607 0.511 805.163 1.266 0.358 4.908 3.377 12.009 LRRNet 1.579 0.473 724.325 1.195 0.358 4.355 3.245 11.529 ALFusion 1.510 0.419 890.521 1.205 0.322 4.878 3.960 9.291 TIMFusion 1.772 0.445 859.554 0.921 0.340 4.249 3.228 11.023 MRFS 1.756 0.514 666.466 1.208 0.415 4.304 3.277 13.071 Text-IF 1.935 0.542 634.406 1.189 0.427 5.277 3.902 11.940 Ours 2.070 0.571 640.661 1.357 0.440 4.144 3.158 14.140
@@
TABLE IV: Quantitative analysis of our method compared with dehazeformer+fusion and text-if on the M3FD dataset. Red indicates the best, and blue indicates the second best.
@@
Methods $Q_{MI}$$\uparrow$ $Q_{AB/F}$$\uparrow$ $Q_{CV}$$\downarrow$ $Q_{SCD}$$\uparrow$ $Q_{VIF}$$\uparrow$ $Q_{NIQE}$$\downarrow$ $Q_{PI}$$\downarrow$ $Q_{SF}$$\uparrow$ MLFusion 1.879 0.400 1023.684 1.133 0.393 4.639 3.232 11.510 U2Fusion 1.647 0.497 824.810 1.281 0.355 5.251 3.635 11.295 LRRNet 1.592 0.463 737.595 1.206 0.360 4.528 3.455 11.116 ALFusion 1.503 0.403 922.165 1.227 0.324 5.095 4.196 8.974 TIMFusion 1.719 0.449 814.112 0.919 0.336 4.420 3.357 11.325 MRFS 1.788 0.509 693.528 1.200 0.417 4.467 3.460 12.551 Text-IF 1.935 0.542 634.406 1.189 0.427 5.277 3.902 11.940 Ours 2.070 0.571 640.661 1.357 0.440 4.144 3.158 14.140
@@
Figure 5: Visual comparison of fusion results from different methods on the RoadScene dataset. In the fusion results generated by the comparison methods, except for the last column, the first row of each image pair represents the result of first dehazing with DIACMP and then fusing. The second row shows the result of dehazing with Dehazeformer and then fusing. The last column represents the result of fusion using Text IF (first row) and our method (second row).
@@
TABLE V: Quantitative analysis of our method compared with diacmp+fusion and text-if on the RoadScene dataset. Red indicates the best, and blue indicates the second best.
@@
Methods $Q_{MI}$$\uparrow$ $Q_{AB/F}$$\uparrow$ $Q_{CV}$$\downarrow$ $Q_{SCD}$$\uparrow$ $Q_{VIF}$$\uparrow$ $Q_{NIQE}$$\downarrow$ $Q_{PI}$$\downarrow$ $Q_{SF}$$\uparrow$ MLFusion 2.350 0.457 509.388 1.248 0.411 3.703 3.196 12.083 U2Fusion 1.852 0.492 842.539 1.201 0.342 3.856 2.960 12.650 LRRNet 1.947 0.363 622.880 0.827 0.359 3.669 3.009 12.498 ALFusion 1.753 0.327 831.380 0.956 0.273 4.311 4.191 8.793 TIMFusion 2.349 0.356 699.810 0.881 0.425 4.484 4.241 10.307 MRFS 2.115 0.391 477.748 1.362 0.393 3.932 3.615 11.269 Text-IF 2.256 0.583 497.001 1.456 0.418 3.737 3.049 13.746 Ours 2.673 0.511 454.497 1.399 0.446 3.333 2.806 14.804
@@
TABLE VI: Quantitative analysis of our method compared with dehazeformer+fusion and text-if on the RoadScene dataset. Red indicates the best, and blue indicates the second best.
@@
Methods $Q_{MI}$$\uparrow$ $Q_{AB/F}$$\uparrow$ $Q_{CV}$$\downarrow$ $Q_{SCD}$$\uparrow$ $Q_{VIF}$$\uparrow$ $Q_{NIQE}$$\downarrow$ $Q_{PI}$$\downarrow$ $Q_{SF}$$\uparrow$ MLFusion 2.425 0.463 509.536 1.239 0.411 3.795 3.343 11.364 U2Fusion 1.864 0.495 862.820 1.175 0.340 3.873 3.029 12.117 LRRNet 1.948 0.360 617.051 0.780 0.357 3.534 2.967 12.326 ALFusion 1.726 0.315 844.762 0.934 0.257 4.321 4.258 8.545 TIMFusion 2.359 0.354 694.572 0.869 0.415 4.457 4.268 10.099 MRFS 2.119 0.396 472.542 1.332 0.389 3.847 3.619 10.922 Text-IF 2.256 0.583 497.001 1.456 0.418 3.737 3.049 13.746 Ours 2.673 0.511 454.497 1.399 0.446 3.333 2.806 14.804
@@
Experiments on M3FD dataset. To evaluate the generalization ability of our method on the M3FD dataset, we select four pairs of IR-VIS images, with the visualization results shown in the Fig. 4. As illustrated in the first and second rows of Fig. 4, most methods exhibit blurring on the store signs, whereas our method preserves clear edges and texture details. Although MLFusion can also retain the texture information of the signs effectively, it performs poorly in restoring background details, resulting in background blurriness. This issue is also evident in the second and third image pairs, where MLFusion and other methods display color distortion in the sky regions. Thanks to the incorporation of infrared information during the dehazing stage in our method, the nextwork effectively restores these sky regions, achieving superior restoration and fusion results.
Quantitative comparison results on M3FD test set, based on the DIACMP and Dehazeformer dehazing methods, are presented in Tables III and IV, respectively. It can be observed that our proposed method ranks first in seven evaluation metrics and second in $Q_{CV}$, indicating its superior restoration and fusion capabilities.
Experiments on RoadScene dataset. We select four pairs of IR-VIS images from the RoadScene dataset to further evaluate the effectiveness and generalization capability of our method. The qualitative comparison results are presented in the Fig. 5. As shown in the first set of results in Fig. 5, our approach provides sharper object edges and more detailed textures. In the third and fourth rows, the content on the billboard within the red box is significantly clearer in our method compared to others, where varying degrees of blurriness are observed. Notably, the content displayed by Text-IF is completely unrecognizable.The results in the fifth and sixth rows indicate that our method can generate fused images with high contrast, preserving the original colors of the signboards while maintaining clear edges. From the results in the red box in the seventh and eighth rows, it can be seen that our method clearly highlights the information of distant vehicles, whereas the results generated by MLFusion, MRFS, and Text-IF are relatively blurry. Although other comparison methods can somewhat enhance vehicle information, as shown in the results in the purple box, their ability to restore the distant sky and roof areas is inferior to that of our method.
Quantitative comparison results on the RoadScene dataset are presented in Tables V and VI, which utilize the DIACMP and Dehazeformer dehazing methods respectively. Our method ranks first in six evaluation metrics, with $Q_{AB/F}$ and $Q_{SCD}$ ranking second. The best $Q_{NIQE}$ and $Q_{PI}$ scores indicate that our method generates fused images that are both natural and sharp.

### v3

@@ 4.3 Comparison with State-of-the-art Methods
Additionally, we conduct a quantitative comparison on 361 image pairs from the MSRS dataset to verify the effectiveness of our method. Tables 1 and 2 present the experimental results based on the DIACMP and Dehazeformer dehazing methods, respectively. As shown in Tables 12, our method ranks first across all eight metrics, demonstrating its outstanding performance on the MSRS dataset. The higher $Q_{AB/F}$ and $Q_{CV}$ scores indicate that our fused images achieve superior detail clarity. Meanwhile, the best $Q_{SCD}$ and $Q_{VIF}$ scores suggest that our method ensures better visual consistency, closely matching the clear source images. The $Q_{MI}$ metric reflects the shared information between the fused and source images, confirming that our approach preserves more source image information. As perceptual quality metrics, $Q_{NIQE}$ and $Q_{PI}$ evaluate image naturalness and perceptual quality in a no-reference manner, where lower values indicate better alignment with human visual perception. Furthermore, $Q_{SF}$ measures the sharpness of image edges, demonstrating that our method produces haze-free fused images that are both natural and sharp.
@@
Figure 4: Visual comparison of fusion results from different methods on the MSRS dataset. In the fusion results generated by the comparison methods, except for the last column, the first row of each image pair represents the result of first dehazing with DIACMP and then fusing. The second row shows the result of dehazing with Dehazeformer and then fusing. The last column represents the result of fusion using Text-IF (first row) and our method (second row).
@@
Table 1: Quantitative analysis of our method compared with DIACMP+fusion and text-if on the MSRS dataset. The best and second-best performances are highlighted with Red and Blue backgrounds, respectively.
@@
Methods $Q_{MI}\uparrow$ $Q_{AB/F}\uparrow$ $Q_{CV}\downarrow$ $Q_{SCD}\uparrow$ $Q_{VIF}\uparrow$ $Q_{NIQE}\downarrow$ $Q_{PI}\downarrow$ $Q_{SF}\uparrow$ MLFusion 1.706 0.274 608.391 1.121 0.199 5.022 4.285 9.953 U2Fusion 1.301 0.330 838.330 1.364 0.226 4.977 4.354 7.209 LRRNet 1.864 0.441 614.101 1.000 0.288 4.750 4.333 8.479 ALFusion 1.449 0.377 700.180 1.201 0.257 5.391 5.419 7.417 TIMFusion 1.800 0.382 1048.770 1.090 0.295 4.515 4.256 8.136 MRFS 1.580 0.457 332.068 1.348 0.325 5.068 4.800 9.042 SHIP 2.021 0.491 509.112 1.316 0.338 5.287 4.723 8.512 FreeFusion 1.892 0.512 441.721 1.214 0.287 4.996 4.478 7.783 Text-IF 1.428 0.558 458.252 1.351 0.369 4.260 4.312 8.393 Ours 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050
@@
Table 2: Quantitative analysis of our method compared with dehazeformer+fusion and text-if on the MSRS dataset. The best and second-best performances are highlighted with Red and Blue backgrounds, respectively.
@@
Methods $Q_{MI}\uparrow$ $Q_{AB/F}\uparrow$ $Q_{CV}\downarrow$ $Q_{SCD}\uparrow$ $Q_{VIF}\uparrow$ $Q_{NIQE}\downarrow$ $Q_{PI}\downarrow$ $Q_{SF}\uparrow$ MLFusion 1.720 0.270 613.291 1.119 0.199 5.052 4.337 9.826 U2Fusion 1.311 0.324 832.878 1.361 0.225 5.185 4.439 7.071 LRRNet 1.883 0.436 608.019 0.999 0.288 4.883 4.445 8.317 ALFusion 1.471 0.370 695.816 1.198 0.256 5.468 5.542 7.339 TIMFusion 1.785 0.382 1057.931 1.092 0.299 4.567 4.355 8.303 MRFS 1.603 0.448 328.901 1.340 0.323 5.150 4.905 8.843 SHIP 2.133 0.486 504.067 1.249 0.328 5.266 4.679 8.448 FreeFusion 1.965 0.498 438.963 1.196 0.276 4.985 4.556 7.675 Text-IF 1.428 0.558 458.252 1.351 0.369 4.260 4.312 8.393 Ours 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050
Experiments on M3FD dataset.
@@
Figure 5: Visual comparison of fusion results from different methods on the M3FD dataset. In the fusion results generated by the comparison methods, except for the last column, the first row of each image pair represents the result of first dehazing with DIACMP and then fusing. The second row shows the result of dehazing with Dehazeformer and then fusing. The last column represents the result of fusion using Text-IF (first row) and our method (second row).
To evaluate the generalization ability of our method on the M3FD dataset, we select four pairs of IR-VIS images, with the visualization results shown in the Fig. 5. As illustrated in the first and second rows of Fig. 5, most methods exhibit blurring on the store signs, whereas our method preserves clear edges and texture details. Although MLFusion is able to preserve the texture information of the signs to some extent, it performs poorly in restoring background details, resulting in noticeable blurring in the background regions. This issue can be clearly observed in the building areas on both the left and right sides of the first image pair. In these regions, the results produced by MLFusion appear overly smooth, with reduced contrast in wall surfaces and window-frame details. This issue is also evident in the second and third image pairs, where MLFusion and other methods display color distortion in the sky regions. Thanks to the incorporation of infrared information during the dehazing stage in our method, the nextwork effectively restores these sky regions, achieving superior restoration and fusion results.
@@
Table 3: Quantitative analysis of our method compared with diacmp+fusion and text-if on the M3FD dataset. The best and second-best performances are highlighted with Red and Blue backgrounds, respectively.
@@
Methods $Q_{MI}\uparrow$ $Q_{AB/F}\uparrow$ $Q_{CV}\downarrow$ $Q_{SCD}\uparrow$ $Q_{VIF}\uparrow$ $Q_{NIQE}\downarrow$ $Q_{PI}\downarrow$ $Q_{SF}\uparrow$ MLFusion 1.842 0.411 937.069 1.134 0.390 4.840 3.230 12.025 U2Fusion 1.607 0.511 805.163 1.266 0.358 4.908 3.377 12.009 LRRNet 1.579 0.473 724.325 1.195 0.358 4.355 3.245 11.529 ALFusion 1.510 0.419 890.521 1.205 0.322 4.878 3.960 9.291 TIMFusion 1.772 0.445 859.554 0.921 0.340 4.249 3.228 11.023 MRFS 1.756 0.514 666.466 1.208 0.415 4.304 3.277 13.071 SHIP 1.929 0.529 676.327 1.258 0.424 4.657 3.475 11.920 FreeFusion 1.865 0.538 986.925 1.272 0.361 4.437 3.235 13.760 Text-IF 1.935 0.542 634.406 1.189 0.427 5.277 3.902 11.940 Ours 2.070 0.571 640.661 1.357 0.440 4.144 3.158 14.140
@@
Table 4: Quantitative analysis of our method compared with dehazeformer+fusion and text-if on the M3FD dataset. The best and second-best performances are highlighted with Red and Blue backgrounds, respectively.
@@
Methods $Q_{MI}\uparrow$ $Q_{AB/F}\uparrow$ $Q_{CV}\downarrow$ $Q_{SCD}\uparrow$ $Q_{VIF}\uparrow$ $Q_{NIQE}\downarrow$ $Q_{PI}\downarrow$ $Q_{SF}\uparrow$ MLFusion 1.879 0.400 1023.684 1.133 0.393 4.639 3.232 11.510 U2Fusion 1.647 0.497 824.810 1.281 0.355 5.251 3.635 11.295 LRRNet 1.592 0.463 737.595 1.206 0.360 4.528 3.455 11.116 ALFusion 1.503 0.403 922.165 1.227 0.324 5.095 4.196 8.974 TIMFusion 1.719 0.449 814.112 0.919 0.336 4.420 3.357 11.325 MRFS 1.788 0.509 693.528 1.200 0.417 4.467 3.460 12.551 SHIP 1.908 0.525 687.446 1.239 0.423 4.779 3.578 11.720 FreeFusion 1.860 0.533 990.661 1.233 0.358 4.530 3.339 13.640 Text-IF 1.935 0.542 634.406 1.189 0.427 5.277 3.902 11.940 Ours 2.070 0.571 640.661 1.357 0.440 4.144 3.158 14.140
Quantitative comparison results on M3FD test set, based on the DIACMP and Dehazeformer dehazing methods, are presented in Tables 3 and 4, respectively. It can be observed that our proposed method ranks first in seven evaluation metrics and second in $Q_{CV}$, indicating its superior restoration and fusion capabilities.
Experiments on RoadScene dataset.
@@
Figure 6: Visual comparison of fusion results from different methods on the RoadScene dataset. In the fusion results generated by the comparison methods, except for the last column, the first row of each image pair represents the result of first dehazing with DIACMP and then fusing. The second row shows the result of dehazing with Dehazeformer and then fusing. The last column represents the result of fusion using Text-IF (first row) and our method (second row).
We select four pairs of IR-VIS images from the RoadScene dataset to further evaluate the effectiveness and generalization capability of our method. The qualitative comparison results are presented in the Fig. 6. As shown in the first set of results in Fig. 6, our approach provides sharper object edges and more detailed textures. In the third and fourth rows, the content on the billboard within the red box is significantly clearer in our method compared to others, where varying degrees of blurriness are observed. Notably, the content displayed by Text-IF is completely unrecognizable.The results in the fifth and sixth rows indicate that our method can generate fused images with high contrast, preserving the original colors of the signboards while maintaining clear edges. From the results in the red box in the seventh and eighth rows, it can be seen that our method clearly highlights the information of distant vehicles, whereas the results generated by MLFusion, MRFS, and Text-IF are relatively blurry. Although other comparison methods can somewhat enhance vehicle information, as shown in the results in the purple box, their ability to restore the distant sky and roof areas is inferior to that of our method.
@@
Table 5: Quantitative analysis of our method compared with diacmp+fusion and text-if on the RoadScene dataset. The best and second-best performances are highlighted with Red and Blue backgrounds, respectively.
@@
Methods $Q_{MI}\uparrow$ $Q_{AB/F}\uparrow$ $Q_{CV}\downarrow$ $Q_{SCD}\uparrow$ $Q_{VIF}\uparrow$ $Q_{NIQE}\downarrow$ $Q_{PI}\downarrow$ $Q_{SF}\uparrow$ MLFusion 2.350 0.457 509.388 1.248 0.411 3.703 3.196 12.083 U2Fusion 1.852 0.492 842.539 1.201 0.342 3.856 2.960 12.650 LRRNet 1.947 0.363 622.880 0.827 0.359 3.669 3.009 12.498 ALFusion 1.753 0.327 831.380 0.956 0.273 4.311 4.191 8.793 TIMFusion 2.349 0.356 699.810 0.881 0.425 4.484 4.241 10.307 MRFS 2.115 0.391 477.748 1.362 0.393 3.932 3.615 11.269 SHIP 2.309 0.495 549.782 1.298 0.435 3.695 3.644 13.346 FreeFusion 2.278 0.449 500.116 1.294 0.401 4.552 3.278 12.847 Text-IF 2.256 0.583 497.001 1.456 0.418 3.737 3.049 13.746 Ours 2.673 0.511 454.497 1.399 0.446 3.333 2.806 14.804
@@
Table 6: Quantitative analysis of our method compared with dehazeformer+fusion and text-if on the RoadScene dataset. The best and second-best performances are highlighted with Red and Blue backgrounds, respectively.
@@
Methods $Q_{MI}\uparrow$ $Q_{AB/F}\uparrow$ $Q_{CV}\downarrow$ $Q_{SCD}\uparrow$ $Q_{VIF}\uparrow$ $Q_{NIQE}\downarrow$ $Q_{PI}\downarrow$ $Q_{SF}\uparrow$ MLFusion 2.425 0.463 509.536 1.239 0.411 3.795 3.343 11.364 U2Fusion 1.864 0.495 862.820 1.175 0.340 3.873 3.029 12.117 LRRNet 1.948 0.360 617.051 0.780 0.357 3.534 2.967 12.326 ALFusion 1.726 0.315 844.762 0.934 0.257 4.321 4.258 8.545 TIMFusion 2.359 0.354 694.572 0.869 0.415 4.457 4.268 10.099 MRFS 2.119 0.396 472.542 1.332 0.389 3.847 3.619 10.922 SHIP 2.237 0.461 539.668 1.286 0.426 3.455 3.646 13.238 FreeFusion 2.168 0.427 484.662 1.281 0.392 4.458 3.169 12.647 Text-IF 2.256 0.583 497.001 1.456 0.418 3.737 3.049 13.746 Ours 2.673 0.511 454.497 1.399 0.446 3.333 2.806 14.804
Quantitative comparison results on the RoadScene dataset are presented in Tables 5 and 6, which utilize the DIACMP and Dehazeformer dehazing methods respectively. Our method ranks first in six evaluation metrics, with $Q_{AB/F}$ and $Q_{SCD}$ ranking second. The best $Q_{NIQE}$ and $Q_{PI}$ scores indicate that our method generates fused images that are both natural and sharp.